Name three(3) sets of Pythagorean triplets and prove that they are indeed Pythagorean triplets.
Answer by mananth(16946)

There are many many triplets
PYthagoras theorem states

the sum of squares of the adjacent sides = hypotenuse^2
leg1^2+leg2^2=hypotenuse^2
In a right triangle the largest side is aways the hypotenuse.
First the basic one
3,4,5
3^2+4^2=5^2
(+16=25
25=25
----
5,12,13
5^2+12^2=13^2
25+144=169
169=169
-------------
Third one you can find